Let YY represent the profit (or loss) for a certain company XX years after 1965. Based on the data shown below, a statistician calculates a linear model Y=−2.09X+22.14Y=-2.09X+22.14.

x y
1 20.15
2 17.98
3 16.21
4 12.94
5 11.77
6 9.9



Use the model to estimate the profit in `1967

y =
Answer #1

For 1967:

x = 1967 - 1965 = 2

Hence,

Estimated profit in 1967 = -2.09 * 2 + 22.14 = 17.96